OpenStudy (anonymous):

write 6m/15m in simplest form 2m/15m 3/5 2/5 3m/5m

OpenStudy (anonymous):

both the numberator and denominator have a common factor 2/5

OpenStudy (earthcitizen):

it's numerator actually, yes it 2/5

OpenStudy (anonymous):

6/15 = 2/5 * 3/3 = 2/5 * 1 = 2/5